# Multi: Multimodal Understanding Leaderboard with Text and Images

### v2

@@ 4.6 Evaluation on Multi-Elite
InternVL 20.7 24.8 23.2 0.0 4.8 17.9 21.0 28.8
Yi-VL 26.2 33.0 29.0 8.7 3.2 32.5 22.7 25.4
Gemini Vision 12.4 5.3 21.2 5.8 12.0 6.8 12.0 37.3
@@
Table 6: Performance of models on Multi-Elite.
We conduct evaluations on Multi-Elite, as outlined in Table 6, which includes 500 specifically chosen questions. These questions are selected based on pre-annotated quality and difficulty scores, in addition to the evaluation results on Multi discussed in § 4.4. The selection aims to ensure a distribution that mirrors Multi’s but also brings challenge to strong MLLMs. Yi-VL achieves the highest score on Multi-Elite with 26.2%, while scores for other models vary between 10.5% and 20.7%. This highlights the substantial challenge presented by Multi-Elite, indicating significant potential for improvement in tackling extremely difficult questions that require in-depth image understanding and intricate reasoning across modalities. It is important to highlight the accuracy of multiple answers choosing (MA Acc.) as the most demanding task for MLLMs, necessitating a thorough grasp of the relationships between the choices and the questions, and reflecting model reliability of selecting all answers correctly.
4.7 Evaluation on Multi-Elite with Multi-Extend
Model window size w/o. kn w. kn
InternVL 768 tokens 20.7 19.9 (-0.8)
Yi-VL 4,096 tokens 26.2 21.4 (-4.8)
Qwen-VL 8,192 tokens 10.5 13.0 (+2.5)
Gemini Vision 30,720 tokens 12.4 17.0 (+4.6)
GPT-4V 128,000 tokens 14.0 21.3 (+7.3)
Table 7: Performance of models with Multi-Extend on Multi-Elite.
@@ 4.7 Evaluation on Multi-Elite with Multi-Extend
The significant challenges posed by Multi-Elite prompt further investigation into the In-Context Learning (ICL) capabilities of MLLMs through the utilization of the Multi-Extend knowledge set. This set is designed to include relevant concepts and frequently utilized solutions related to the problems. The study is conducted on several MLLMs, with the prompts for incorporating these knowledge pieces shown in Figure 5, and the results are listed in Table 7. Notably, the average number of tokens per question escalates from 65 to 250, and further to 850, following the integration of prompts and the adoption of Multi-Extend, with the most extensive examples surpassing 10,000 tokens. Multi-Extend poses a significant challenge in terms of the necessary window size and the capacity to handle lengthy contexts. It is observed that models equipped with larger window sizes, i.e. Gemini Vision and GPT-4V, benefit more from Multi-Extend, whereas there is a notable decline in performance for MLLMs with smaller window sizes. The increase in tokens may also present a hurdle for models, as the concise question may become overshadowed by the extensive context.
@@ 4.8 Takeaways
  • GPT-4V demonstrates the highest performance with a 63.7% score, indicating a significant challenge of Multi, while Yi-VL leads among open-source models.
  • MLLMs show a performance drop in questions requiring more images, with only GPT-4V exceeding a basic guessing baseline in multi-image scenarios.
  • LLMs show a reverse correlation in performance between non-image and single/multiple image sets, highlighting the challenge of avoiding hallucination in visual questions.
  • Models generally perform better on questions requiring shorter answers, i.e. SA > MA > FB > OP. The results of MA Acc. emphasize the importance of balancing recall and precision.
  • Performance trends are consistent across educational levels, with lower scores on AAT questions due to their multimodal complexity.
  • The inclusion of images significantly boosts question-answering performance, with captions offering a slight improvement and OCR text potentially complicating reasoning processes.
  • In the Multi-Elite evaluation, Yi-VL achieves the highest 26.2% score, illustrating the difficulty of Multi-Elite and the need for advanced image understanding and reasoning across modalities.
  • The aid of Multi-Extend help improve performance on models with long window sizes, yet it may yield adverse effects on less capable models.
@@ 5 Conclusion
In this paper, we introduce Multi, a comprehensive and challenging benchmark designed to rigorously evaluate the performance of MLLMs in detailed cross-modality understanding and scientific reasoning. Our experiments with state-of-the-art models like Qwen-VL, InternVL, Yi-VL, Gemini, and GPT-4 demonstrate that while these models exhibit promising capabilities, there remains a significant gap compared to human performance, particularly in tasks involving cross-modal alignment, logical reasoning, and complex comprehension. This underscores the need for continuous research and development in this domain.
@@
Limitations and Future Work
Multilingual Capabilities
@@
Multi predominantly features simplified Chinese and mainly focuses on subjects taught in Chinese schools, with limited English multimodal content that’s relatively straightforward for LLMs. We plan to include translations in future versions. Nonetheless, the presence of Chinese characters in figures poses a significant challenge for MLLMs trained on different linguistic datasets.
Use of Explanations
@@
While we have annotated explanations in detail, the utilization in subsequent studies remains limited. These explanations could potentially serve as valuable training data for model fine-tuning and few-shot learning using methods like CoT (Chain-of-Thoughts) or RAG (Retrieval Augmented Generation) and may aid in evaluating reasoning skills.
Metrics for evaluating blank-filling, open-ended writing and others
@@
Our evaluation primarily uses exact match, which might be overly stringent for assessing MLLMs’ true capabilities. Assessing open-ended writing tasks that require complex knowledge and reasoning is still a challenge. We also have 100 questions that do not belong to traditional categories, such as questions requiring geographic drawing, and the evaluation of them will be even more challenging. Now that only a few studies Wang et al. (2023) involve human evaluation, developing automatic and reliable methods remains an open research area.
Adaptation to various MLLMs
@@
Although we have tested several MLLMs, numerous others exist and new ones are continuously emerging. We encourage the community to evaluate their MLLMs using our benchmark to gauge their cognitive reasoning abilities. We will test more models as soon as the multilingual version is released.
Expansion to more modalities and subjects
Our benchmark currently focuses on static images, but incorporating other modalities like audio and video, and subjects like art, music theory, medicine, and sports could present new topics. Thus, expanding our question set to cover these areas is a promising direction for future research.
Appendix A Statistics
We providedetailed statistics in Table 8. One question may contain more than one scoring points as mentioned in § 4.3.
Statistics Number Points
Total Problems 17251 -
Total Questions 18430 -
Total Points 23320 -
Total Images 7658 -
Total Knowledge 4595 -
Multiple Choices 16100(87.36%) 19904(85.35%)
 - Single Answer 13963(75.76%) 13963(59.88%)
 - Multiple Answers 2137(11.60%) 5941(25.48%)
Fill in the Blank 1432(7.77%) 2211(9.48%)
Open Ended Writing 798(4.33%) 1205(5.17%)
Others 100(0.54%) -
Question with Images 7489(40.63%) 9042(38.77%)
 - Single Image 7265(39.42%) 8767(37.59%)
 - Choices within Image 1179(6.40%) 1181(5.06%)
 - Multiple Images 224(1.22%) 275(1.18%)
Question with Explanations 10565(57.33%) 13186(56.54%)
Question with Knowledge 9048(49.09%) 12919(55.40%)
Table 8: The statistic overview of Multi.
A.1 Data Distribution on Question Types
Type # choices # A # B # C # D # E,F,G…
SA 2 1819 1376 0 0 0
3 272 287 262 0 0
4 2193 2638 2708 2379 0
5 0 2 7 9 0
MA 3-13 1467 1568 1510 1303 91
Total 2-13 5751 5871 4487 3691 91
Table 9: The choice distribution for multiple-choice questions.
In addition to multiple-choice questions, our benchmark also includes a substantial number of fill-in-the-blank and open-ended questions, creating a diverse and comprehensive range of testing scenarios. Moreover, we have incorporated unique open-response questions that require creative answers, such as drawings. It is important to note that these open-response questions are not included in our formal evaluation and scoring procedures; they are primarily proposed for qualitative research and development in the field of MLLMs. Our benchmark is carefully designed to thoroughly assess and enhance the ability of MLLMs to process and respond to various question types, resembling real-world scenarios.
Appendix B Models

### v3

@@ 4.5 Is Multi-Elite more challenging?
Yi-VL-34B 26.2 33.0 29.0 8.7 3.2 32.5 22.7 25.4
InternVL-1.1 20.7 24.8 23.2 0.0 4.8 17.9 21.0 28.8
InternVL2-8B 36.3 38.5 39.5 27.6 23.3 40.4 39.5 23.3
MiniCPM-V-2.6 21.3 23.3 26.1 7.6 5.8 18.4 22.6 23.4
Qwen2-VL-7B 40.0 43.5 43.3 32.4 23.3 39.9 38.4 50.0
Qwen2-VL-72B 53.1 55.3 62.8 54.3 25.8 63.6 49.6 37.5
@@
Gemini Vision Pro 12.4 5.3 21.2 5.8 12.0 6.8 12.0 37.3
GPT-4o-mini 18.2 22.3 18.5 11.5 7.2 17.1 17.4 28.8
GPT-4o (0806) 36.0 43.1 37.5 24.0 15.2 34.2 35.9 44.1
Gemini 1.5 Flash 33.4 32.3 38.3 21.9 25.8 33.8 30.7 50.0
Claude 3.5 Sonnet 34.3 38.1 38.6 30.8 16.0 33.3 33.3 45.8
GPT-4o (1120) 36.3 42.2 42.1 25.7 17.5 36.0 36.3 56.2
Gemini 1.5 Pro 46.9 47.8 47.9 33.3 42.5 43.4 46.0 65.6
Table 7: The performance of all models evaluated on Multi-Elite. $\dagger$: these models are chosen to guide the Multi-Elite selection process, thus the performance of these models are relatively lower.
Figure 6: Performance improvement with Multi-Extend on Multi-Elite. The context window size of each model is plotted in log scale.
4.6 Can model utilize Multi-Extend?
@@
The significant challenges posed by Multi-Elite prompt further investigation into the In-Context Learning (ICL) capabilities of MLLMs through the utilization of the Multi-Extend knowledge set. This set is designed to include relevant concepts and frequently utilized solutions related to the problems. Notably, the average number of tokens per question escalates from 65 to 250, and further to 850, following the integration of prompts and the adoption of Multi-Extend, with the most extensive examples surpassing 16,000 tokens.
The study is conducted on several selected MLLMs, with the prompts for incorporating these knowledge pieces shown in Figure 10. In Figure 6, we present the performance improvement or decline on Multi-Elite with Multi-Extend provided as external knowledge. We also plot the context window size of each model in the figure.
Multi-Extend poses a significant challenge in terms of the necessary window size and the capacity to handle lengthy contexts. It is observed that models equipped with larger window sizes, i.e. GPT-4 series, are more likely to benifit from Multi-Extend, whereas there is a notable decline in performance for MLLMs with smaller window sizes. The increase in tokens may also present a hurdle for models, as those concise questions may become overshadowed or distracted by the extensive context.
4.7 Do these images provides information?
Model NI SI MI
w/o. w. cap w. ocr w. img w/o. w. cap w. ocr w. img
MOSS 36.1 27.3 27.3 ($+$0.0) 27.6 ($+$0.3) - 17.1 20.7 ($+$3.6) 19.0 ($+$1.9) -
Qwen2 61.6 35.3 31.6 (--3.7) 35.0 (--0.3) - 18.0 16.2 (--1.8) 17.1 (--0.9) -
InternLM2.5 70.3 39.2 38.5 (--0.7) 38.5 (--0.7) - 24.5 28.6 ($+$4.1) 23.3 (--1.2) -
[…]
VisualGLM 35.1 20.8 21.4 ($+$0.6) 20.4 (--0.4) 25.2 ($+$4.4) 15.3 15.1 (--0.2) 14.5 (--0.8) 9.7 (--5.6)
VisCPM 36.8 27.1 27.6 ($+$0.5) 27.2 ($+$0.1) 28.4 ($+$1.3) 24.8 21.6 (--3.2) 20.9 (--3.9) 16.6 (--8.2)
Chinese-LLaVA 32.3 26.1 26.3 ($+$0.2) 25.5 (--0.6) 22.6 (--3.5) 17.6 19.9 ($+$2.3) 19.6 ($+$2.0) 17.8 ($+$0.2)
Qwen-VL 43.2 30.7 30.3 (--0.4) 31.0 ($+$0.3) 32.7 ($+$2.0) 25.5 25.0 (--0.5) 26.2 ($+$0.7) 20.7 (--4.8)
Yi-VL-34B 63.8 39.9 38.7 (--1.2) 39.4 (--0.5) 42.0 ($+$2.1) 24.1 26.5 ($+$2.4) 24.2 ($+$0.1) 24.5 ($+$0.4)
InternVL-1.1 50.9 33.4 33.3 (--0.1) 33.1 (--0.3) 35.5 ($+$2.1) 24.8 21.9 (--2.9) 22.9 (--1.9) 25.1 ($+$0.3)
InternVL2-8B 78.7 48.4 49.0 ($+$0.6) 48.2 (--0.2) 51.6 ($+$3.2) 32.7 32.6 (--0.1) 28.6 (--4.1) 28.7 (--4.0)
MiniCPM-V-2.6 63.8 37.9 35.0 (--2.9) 38.3 ($+$0.4) 41.8 ($+$3.9) 21.8 25.1 ($+$3.3) 24.0 ($+$2.2) 24.0 ($+$2.2)
Qwen2-VL-7B 78.1 47.8 48.9 ($+$1.1) 48.5 ($+$0.7) 53.7 ($+$5.9) 29.2 30.4 ($+$1.2) 30.5 ($+$1.3) 34.2 ($+$5.0)
Qwen2-VL-72B 86.9 52.1 56.6 ($+$4.5) 56.6 ($+$4.5) 61.5 ($+$9.4) 28.4 33.0 ($+$4.6) 32.0 ($+$3.6) 41.1 ($+$12.7)
[…]
Gemini/Vision Pro 62.5 36.2 36.9 ($+$0.7) 38.4 ($+$2.2) 40.0 ($+$3.8) 18.3 23.2 ($+$4.9) 18.6 ($+$0.3) 24.5 ($+$6.2)
ChatGPT 54.0 6.8 9.9 ($+$3.1) 6.6 (--0.2) - 5.1 10.7 ($+$5.6) 5.5 ($+$0.4) -
GPT-4/V 74.5 11.3 9.7 (--1.6) 1.9 (--9.4) 46.9 ($+$35.6) 8.8 9.4 ($+$0.6) 3.1 (--5.7) 28.1 ($+$19.3)
GPT-4o-mini 67.7 17.5 30.0 ($+$12.5) 29.6 ($+$12.1) 43.0 ($+$25.5) 8.2 14.9 ($+$6.7) 10.3 ($+$2.1) 30.4 ($+$22.2)
Gemini 1.5 Flash 73.8 26.3 38.7 ($+$12.4) 38.7 ($+$12.4) 50.9 ($+$24.6) 11.9 21.4 ($+$9.5) 15.5 ($+$3.6) 40.0 ($+$28.1)
GPT-4o (0806) 80.0 18.5 33.6 ($+$15.1) 33.8 ($+$15.3) 54.1 ($+$35.6) 12.0 17.8 ($+$5.8) 13.2 ($+$1.2) 40.9 ($+$28.9)
Gemini 1.5 Pro 80.2 29.4 43.5 ($+$14.1) 47.6 ($+$18.2) 58.5 ($+$29.1) 11.6 27.9 ($+$16.3) 25.4 ($+$13.8) 50.0 ($+$38.4)
[…]
Table 8: Performance with different image information on Multi. w/o.: No image information is provided. w. cap: Image captions are provided. w. ocr: Image OCR transcripts are provided. w. img: Original images from Multi are provided. We calculate the average improvements for each group of models and different image types.
To assess the necessity of images in Multi for solving problems, we conduct an ablation study where we either remove images from the SI and MI sets or substitute them with textual descriptions, such as captions and OCR-derived text. We utilize BLIP2 Li et al. (2023a) for generating image captions and EasyOCR [^4] https://pypi.org/project/easyocr/ to extract text from images. The results are shown in Table 8.
For questions that incorporate a single image (as indicated in the SI column), the presence of images significantly aids in answering the questions, with an average performance boost of 3.08% and 25.13% of open-sourced MLLMs and closed-source MLLMs. Notably, both GPT-4V and GPT-4o (0806) experience a substantial increase of 35.6% in performance, primarily due to their tendency to abstain from answering in the absence of images. We will further discuss this in § 4.8.
For questions that involve multiple images (as discussed in the MI column), we categorize models into three groups: 1) Close-source models, specifically GPT-4V and Gemini Pro, which leverage all images and achieve significant improvement. 2) Open-source models capable of handling multiple high-resolution images at a time, namely Qwen2-VL and MiniCPM-V-2.6, which show promising improvements. 3) Open-source models without multi-image support, like VisualGLM, and VisCPM, which show huge performance decline. Only one of the images given could be distracting, especially in a situation where each choice contains an image.
In settings where images are omitted and replaced by their textual descriptions (captions or OCR text), a small improvement is observed with either captions or OCR text for most models. Captions, which generally summarize the images, introduce bilingual elements to the models and usually miss details. OCR text, while detailed, lacks spatial information and is not universally applicable, as some images contain no text at all. Both forms of textual information lower the models’ refusal rate, and LLMs benefit more from these than MLLMs. However, they potentially complicate reasoning processes. Nevertheless, a generic caption is found to more frequently outperform scattered OCR fragments.
4.8 Will these models refuse to answer if the image is not provided?
Multi not only expects models to achieve higher scores when provided with multimodal information but also aims to investigate the performance and reliability of MLLMs under pure text input conditions. As a result, we observe that some models like the GPT series, which perform significantly better in w. img setting than pure text setting in Table 8.
We further analyzed the outputs of these models as we prompted them to reject if the information given was not sufficient. The details are presented in Figure 7 and are within our expectations. Theoretically, the greater the performance improvement of a model with and without image information, the better it demonstrates superiority in human alignment and control over hallucinations. Strong models like GPT4 and Gemini 1.5 reject a big proportion of questions, indicating the ability to recognize missing information in the question content, which leads to a lower score. These models keep improving with user feedback and are better aligned with human preferences. Conversely, those open-sourced models are less likely to reject such questions and show a non-negligible direction for future improvement.
(a) SI set without image information
(b) MI set without image information
Figure 7: The portion of answer rejection when evaluated on SI set and MI set without image information as input.
4.9 Is Multi sensitive to different prompts?
To investigate the sensitivity of Multi to variations in prompt settings, we conducted an ablation study using multiple prompt configurations. Specifically, we examined the following scenarios: (1) base: the original prompt, (2) eng: using English as the prompt language, (3) no bg: removing the question background in the beginning, i.e. You need to use your knowledge of {knowledge} to answer this {question_type} question, (4) cot: adding zero-shot chain-of-thoughtKojima et al. (2022) reasoning instructions. The results are presented in Table 9.
The results indicate that Multi is stable by switching to another language or drop background information, yet will benefit by chain-of-thoughts reasoning.
Switching to English prompts (eng) resulted in a minor average improvement of 1.0%, suggesting that prompt language has limited but non-negligible influence. This might be due to a higher proportion of English fine-tuning data, which better aligns the models with English instructions. Removing the background knowledge section (no bg) had negligible effects on the results.
Incorporating chain-of-thought reasoning (cot) consistently improves the results for most models, with an average improvement of 2.9%. Notably, GPT-4o-mini showed the most significant benefit from cot, achieving a 12.2% increase. The 8.9% sharp drop in performance for Qwen2-VL-7B can be attributed to its failure to follow to the chain-of-thought prompt, where it directly provided predictions without intermediate reasoning steps in approximately 40% of the cases. This suggests a limitation in the model’s alignment with instruction-following for reasoning-intensive tasks. This trend is also observed in Qwen-VL-72B where 5% of the questions are answered directly.
These findings demonstrate the critical role of well-crafted prompts in optimizing MLLM performance on Multi. However, using different prompt settings does not affect the overall ranking of the models in the experiments for most cases.
Model base eng no bg cot
InternVL2-8B 36.3 35.6 (--0.7) 35.6 (--0.7) 38.3 ($+$2.0)
MiniCPM-V-2.6 21.3 25.0 ($+$3.7) 21.6 ($+$0.3) 26.3 ($+$5.0)
Qwen2-VL-7B 40.0 40.8 ($+$0.8) 39.0 (--1.0) 31.1 (--8.9)
Qwen2-VL-72B 53.1 53.9 ($+$0.8) 54.6 ($+$1.5) 51.1 (--2.0)
GPT-4o-mini 18.5 20.2 ($+$1.7) 21.2 ($+$2.7) 30.7 ($+$12.2)
GPT-4o (1120) 36.3 37.6 ($+$1.3) 38.8 ($+$2.5) 42.0 ($+$5.7)
Gemini 1.5 Flash 33.4 34.0 ($+$0.6) 33.1 (--0.3) 36.4 ($+$3.0)
Gemini 1.5 Pro 46.9 46.1 (--0.8) 45.0 (--1.9) 52.6 ($+$5.7)
[…]
Table 9: The comparison of different prompt settings. base: the original prompt. cot: zero-shot chain-of-thought. eng: English translation of the original prompt. no bg: remove background information.
Figure 8: The required ability distribution of Multi-Elite.
Figure 9: The error distributions of several models’ responses.
4.10 What is the main cause of the error?
To explore the primary causes of errors, we analyzed 165 questions (one-third of Multi-Elite), using predictions generated by the 8 models evaluated in § 4.9. To get more detailed insights, we focused on outputs generated with the CoT prompt, which helps reveal patterns and challenges across different error types. The detailed distribution of all errors is shown in Figure 9, and we give five example cases with error analysis for several selected models as shown in F. We also asked human experts to identify key abilities required to solve each question, such as comprehension of text and images, mathematical calculation, logical reasoning, knowledge recall, and others. The distribution of these abilities is shown in Figure 8, allowing us to compare the alignment between the skills demanded by the questions and the following observed error patterns.
We summarize these errors into five primary categories, with the following observations:
Alignment Errors (11.1%): A significant proportion of errors cause by models failing to follow CoT (Qwen2-VL-7B in Figure 15 and Figure 18) or formatting instructions (InternVL2-8B in Figure 18), highlighting difficulties in instruction compliance. Some models also rejected questions outright, potentially due to divergences in human alignment objectives. Notably, Qwen2-VL-7B and 72B exhibited higher rejection rates when prompted with CoT but proceeded to provide direct answers, as discussed in § 4.9.
Text Understanding Errors (4.4%): Failures in text comprehension account for 4% to 6% of errors, which is lower than the importance of text understanding identified by human experts. Divided into summary-level and detail-level errors, these failures typically involved an inability to extract critical context words, even though the models performed well on general understanding. Or, conversely, the model pays too much attention to detail and ignores the overall theme (GPT-4o-mini in Figure 19).
Image Understanding Errors (24.4%): Much different from text understanding, errors in image summarization and detailed analysis account for 15% to 30% of total errors. Since 107 (65%) out of the 165 questions analyzed involve images, it demonstrates the challenge of handling image inputs. Similar to text errors, detail-level image comprehension caused more errors, as it is the bottleneck for all models to handle the AAT questions just like the one in Figure 15. There are cases in which models often struggle to locate and focus on specific image elements as mentioned in the text (GPT-4o in Figure 18), sometimes resulting in visual hallucinations. Enhanced strategies for analyzing and synthesizing image details are crucial to address this issue.
Reasoning Errors (21.7%): Logical reasoning errors emerged primarily in tasks requiring multi-step deductions (MiniCPM-V-2.6 in Figure 16) or calculations after extracting relevant information from text and image contexts. There is also the possibility that the model just repeats its ineffective reasoning steps for all options (InternVL2-8B in Figure 19). These errors underscore the difficulty of complex reasoning, which remains a significant bottleneck for both LLMs and MLLMs.
Lack of Knowledge or Common Sense (35.8%): Errors not attributed to the categories above often resulted from insufficient domain-specific knowledge. Models frequently failed to retrieve or apply relevant knowledge to solve or even comprehend the questions (GPT-4o-mini in Figure 15 and Figure 17). As discussed in § 4.6, adding domain-specific knowledge has shown improvement for certain models, provided they can utilize the added information without being misled. Yet there is much to learn for models in training stages, and a challenge to effectively adopt retrieval augmented generation (RAG).
System Bias (2.6%): A small proportion of errors resulted from annotation inaccuracies or answer extraction issues, categorized as system bias. These errors have been addressed and corrected in the latest dataset release.
In summary, the observed error types align closely with the skill requirements identified in the dataset, reflecting the persistent challenges in reasoning and multimodal comprehension for MLLMs, even with CoT prompting. Addressing these obstacles will require significant advancements in model alignment, instruction following, and cross-modal reasoning capabilities.
@@ 4.11 Takeaways
  • Qwen2-VL-72B demonstrates the highest performance with a 76.9% score, indicating a significant challenge of Multi, while Gemini 1.5 Pro leads among closed-source models with 71.8%. There is still a large gap between most capable MLLMs and human experts in general, while in some subjects like the Driving Test, Qwen2-VL-72B has surpassed the human baseline, but the AAT test still remains as the greatest challenge for all models. (Table 4 in § 4.4)
  • MLLMs show a performance drop in questions requiring more images, with only Gemini 1.5 Pro reaching 50.0% of the score rate in multi-image scenarios. (Table 5 in § 4.4)
  • Models generally perform better on questions requiring shorter answers, i.e. SA $>$ MA $>$ FB $>$ OP. The results of MA Acc. emphasize the importance of balancing recall and precision, where less capable MLLMs failed to provide all correct choices. (Table 6 in § 4.4)
  • In the Multi-Elite evaluation, the gap between models and human experts becomes greater by 20.0%, illustrating the difficulty of Multi-Elite and the need for advanced image understanding and reasoning across modalities. (Table 7 in § 4.5)
  • The inclusion of Multi-Extend is likely to help improve performance on models with longer window sizes. However, it might also have adverse effects due to multiple reasons. (Figure 6 in § 4.6)
  • Including image information in questions significantly enhances model performance, especially when original images are included, though textual replacements like captions or OCR text provide some utility but lack comprehensive spatial and detail fidelity. (Table 8 in § 4.7)
  • Strong models like GPT-4 series and Gemini 1.5 effectively refuse to answer when critical image information is missing, showcasing better alignment with human preferences, while open-source models are likely to answer based on available text information. (Figure 7 in § 4.8)
  • Prompt variations affect performance, with chain-of-thought reasoning enhancing accuracy while removing system-level instructions leads to notable declines. The relative performance across models remains stable and robust to different prompt settings. (Table 9 in § 4.9)
  • Errors in model performance are primarily driven by failures in knowledge recall, logical reasoning, and image understanding. Additionally, alignment and instruction-following issues underscore challenges in maintaining adherence to complex prompts, particularly in chain-of-thought tasks. The observed error types match closely with the skill requirements identified by human annotators. (Figure 8 and Figure 9 in § 4.10)
@@ 5 Conclusion
In this paper, we introduce Multi, a comprehensive and challenging benchmark designed to rigorously evaluate the performance of MLLMs in detailed cross-modality understanding and scientific reasoning. Our experiments with state-of-the-art models like Qwen2-VL, InternVL2, Gemini 1.5, and GPT-4 series demonstrate that while these models exhibit promising capabilities, there remains a significant gap compared to human performance, particularly in tasks involving cross-modal alignment, logical reasoning, and complex comprehension. This underscores the need for continuous research and development in this domain.
@@
Appendix A Limitations and Future Work
  • Potential biases: While we have taken several steps to minimize biases in the dataset and ensure fair evaluation , we acknowledge that some biases remain. Despite our efforts, challenges in the construction and evaluation process, such as subject, difficulty, format, and system biases, still persist to some extent. The distribution of questions across different subjects and difficulty levels may still exhibit some inherent imbalances, which could impact the evaluation of MLLMs in specific domains or tasks. Although we have adopted algorithms for subject selection and multi-round annotation for difficulty balancing, these processes are not foolproof. Additionally, variations in the annotation process, such as subjective difficulty ratings and semantic rephrasing of questions, may introduce slight inconsistencies across different subsets of the data. We are committed to refining our methodology in future versions to further mitigate these biases and enhance the fairness of the benchmark.
@@ Appendix A Limitations and Future Work
  • Multilingual capabilities: Multi predominantly features simplified Chinese and mainly focuses on subjects taught in Chinese schools, with limited English multimodal content that’s relatively straightforward for MLLMs. We plan to include translations in future versions. Nonetheless, the presence of Chinese characters in figures poses a significant challenge for MLLMs trained on different linguistic datasets.
@@
  • Use of explanations: While we have annotated explanations in detail, the utilization in subsequent studies remains limited. These explanations could potentially serve as valuable training data for model fine-tuning and few-shot learning using methods like CoT (Chain-of-Thoughts) or RAG (Retrieval Augmented Generation) and may aid in evaluating reasoning skills.
@@
  • Metrics for evaluating blank-filling, open-ended writing and others: Our evaluation primarily uses exact match, which might be overly stringent for assessing MLLMs’ true capabilities. Assessing open-ended writing tasks that require complex knowledge and reasoning is still a challenge. We also have nearly 100 questions that do not belong to traditional categories, such as questions requiring geographic drawing, and the evaluation of them will be even more challenging. Now that only a few studies Wang et al. (2023) involve human evaluation, developing automatic and reliable methods remains an open research area.
@@
  • Adaptation to various MLLMs: Although we have tested several MLLMs, numerous others exist and new ones are continuously emerging. We encourage the community to evaluate their MLLMs using our benchmark to gauge their cognitive reasoning abilities. We will test more models as soon as the multilingual version is released.
Appendix B Models and Experiment Cost